# Spin-Dependent Force and Inverted Harmonic Potential for Rapid Creation of Macroscopic Quantum Superpositions

### v1

@@ Abstract
Creating macroscopic spatial superposition states is crucial for investigating matter-wave interferometry and advancing quantum sensor technology. Currently, two potential methods exist to achieve this objective. The first involves using inverted harmonic potential (IHP) to spatially delocalize quantum states through coherent inflation [1]. The second method employs a spin-dependent force to separate two massive wave packets spatially [2]. The disadvantage of the former method is the slow initial coherent inflation, while the latter is hindered by the diamagnetism of spin-embedded nanocrystals, which suppresses spatial separation. In this study, we integrate two methods: first, we use the spin-dependent force to generate initial spatial separation, and second, we use IHP to achieve coherent inflating trajectories of the wavepackets. This approach enables the attainment of massive large spatial superposition in minimal time. For instance, a spatial superposition with a mass of $10^{-15}$ kg and a size of 50 $\mu$m is realized in $0.1$ seconds. We also calculate the evolution of wave packets in both harmonic potential (HP) and IHP using path integral approach.
@@ I Introduction
The significant interest in massive quantum superposition states primarily arises from three aspects. The first is the exploration of the quantum-classical boundary [3]. Quantum interference, from electrons to macromolecules ($10^{-31}-10^{-23}$ kg), has been observed in contemporary experiments, demonstrating their quantum nature [4, 5, 6, 7, 8]. This leads to quests about whether quantum superposition states can also be achieved for objects of larger mass [9]. The second aspect is their utility in validating theoretical models [10]. For instance, they can be employed to test wave function collapse theories [11, 12], modified quantum mechanical frameworks [13, 14, 15], and examine the weak equivalence principle [16, 17]. Additionally, they may reveal the quantum nature of gravity by combining two massive spatial superposition states [18, 19, 20, 21], see theoretical explanations  [22, 23, 24, 25, 26, 27, 28].
The third aspect is that they can act as highly sensitive quantum sensors, detecting phenomena such as the Casimir force and dipole interactions [29, 30, 31, 32], gravitational waves [33, 34], quantum sensors for detecting accelerations, and inertial rotations [35, 36, 37], dark matter [38], physics beyond the Standard Model [39], testing massive graviton [40], non-local gravitational interaction [22, 41], and analogue of light bending experiment in quantum gravity [42].
@@
Based on these cutting-edge techniques, numerous experimental schemes for creating macroscopic spatial superposition states have been proposed [15, 2, 54, 55, 1, 56, 57, 58, 59, 60, 32]. A natural method to achieve a delocalized quantum state is to allow a quantum wave packet to evolve freely [15]. However, this delocalization process is slow. For instance, consider a silica microsphere with a mass of $10^{-15}$ kg trapped by a magnetic field at a frequency of 100 Hz and cooled to its ground state [61]. Its initial wave packet spatial width is approximately $10^{-11}$ m. After 1 s of free evolution, the wave packet width becomes about $10^{-9}$ m, roughly one-thousandth of its size. To accelerate this delocalization process, an IHP can induce coherent inflation [1, 62]. This approach allows the coherent length of a $10^{-14}$ kg nanoparticle to increase to around 1 $\mu$m in 0.6 s, making the coherence length comparable to its size. Another method for creating macroscopic spatial superposition states involves utilizing spin-dependent forces, such as using diamond embedded with a NV center [2, 54, 57, 56]. Initially, a pulse is used to place the electron spin of the NV center in a superposition state, followed by applying a magnetic field to induce spatial splitting, similar to the Stern-Gerlach experiment. The advantage of this method is the ease of preparing the initial superposition state and reading out the final spin state [63]. However, the diamagnetism of the diamond suppresses the spatial separation of the wave packets when an external magnetic field is applied [59, 64, 65, 66], limiting the direct increase of the superposition size through enhanced magnetic field gradients.
@@
The paper is organized as follows. Section II presents the specific experimental protocol and the magnetic fields required to construct the HP and IHP for the experiment. Section III provides an analytical solution for the classical trajectories of the wave packets at each experiment stage. In Section IV, we numerically calculate the classical trajectories of the wave packet without approximations for the nonlinear magnetic field and compare these results with the analytical solution. Section V discusses the evolution of the wave packet under HP and IHP using path integrals. In Section VI, we examine the effect of fluctuations in the magnetic field gradient on wave packet contrast for both the HP and IHP cases. Finally, we conclude our findings in Section VII.
@@ II Experimental scheme
$\hat{\mathbf{H}}=\frac{1}{2m}\hat{\mathbf{P}}^{2}-\frac{\chi_{\rho}m}{2\mu_{0}}\hat{\mathbf{B}}^{2}+\hbar\gamma_{e}\hat{\mathbf{S}}\dotproduct\hat{\mathbf{B}}+\hbar D\hat{S}_{z}^{2},$ (1)
@@
The first term represents the kinetic energy of the nanodiamond, where $\hat{\mathbf{P}}$ is the momentum and $m$ is the mass of the nanodiamond. The second term signifies the magnetic energy of a diamagnetic material (nanodiamond) in a magnetic field, with $\chi_{\rho}=-6.2\times 10^{-9}\,\,\text{m}^{3}/\text{kg}$ as the mass susceptibility and $\mu_{0}$ as the vacuum permeability. The third term is electron spin interacting with the magnetic field. $\hbar$ is the reduced Plank constant and $\gamma_{e}$ the electronic gyromagnetic ratio. $\hat{\mathbf{S}}$ is the spin operator. $\hat{\mathbf{B}}$ is the magnetic field. The last term represents the zero-field splitting of the NV center with $D=(2\pi)\times 2.8$ GHz. $\hat{S}_{z}$ is the spin component operator aligned with the NV axis. In this paper, we will assume that the nanodiamond’s rotational angular momentum is zero; see for a finite angular momentum and its implications [64]
The experimental protocol is divided into five stages: the initial separation stage, the enhancement stage, the return stage, the deceleration stage, and the recombination stage. This division aims to achieve a large spatial superposition size in a short time (around 0.1 s) and to recombine the trajectories of the CoM, thereby restoring spin coherence [51]. The five stages are outlined in Fig. 1 and described below.
@@
Figure 1: The sketch of the experimental scheme. The light orange quadratic curves represent harmonic or IHPs. The blue and orange curves with shaded fills represent wave packets in a superposition state. The arrows inside the wave packet represent the corresponding electron spin state (spin up or spin down). The arrows along the potential indicate the direction of motion of the wave packet. The double arrows in the return stage indicate that the wave packet first separates and then returns. The time axis from top to bottom represents the initial separation stage, the enhancement stage, the return stage, the deceleration stage, and the recombination stage, or stages 1-5, respectively. $T_{1}$ to $T_{5}$ denote the evolutionary time of each stage. The grey dashed lines correspond to the moments at which the $\pi/2$ pulse is applied.
  • Initial state: The initial state is a spin superposition state (|$(\ket{+1}+\ket{-1})/\sqrt{2}$, and |$\ket{+1}$ and |$\ket{-1}$ are eigenstates with eigenvalues of 1 and -1 for the spin operator $\hat{\mathbf{S}}$.
  • Initial separation stage (Stage 1): Switching on an external linear magnetic field, the wave packets with different spin eigenstates generate spatial separation in the HP.
  • Enhancement stage (Stage 2): When the spatial separation of the two wave packets in the HP is maximized, the linear magnetic field is switched off. Then a $\pi/2$ pulse is applied to transform the state (|$(\ket{+1}+\ket{-1})/\sqrt{2}$ into |$\ket{0}$ [68, 69]. We ignore the time of applying the pulse here. Subsequently, switching on the nonlinear magnetic field creates an IHP allowing the wave packets to separate rapidly, increasing the superposition size in a short period of time.
  • Return stage (Stage 3): The nonlinear magnetic field is switched off, and the linear field is switched on after the wave packet has evolved in the IHP for a while  [^1] The exact evolution time is related to the superposition size we want to achieve. The longer the time, the larger the spatial separation between the wave packets.. The velocities of the wave packets gradually decrease in the HP and then reverse. The spatial separation between wave packets reaches its maximum at this stage.
  • Deceleration stage (Stage 4): When the magnitude of the wave packet velocity is the same as the magnitude at the end of the second stage and in the opposite direction, the linear magnetic field is switched off and the nonlinear magnetic field is switched on. The wave packet gradually decelerates in the IHP.
@@
  • Recombination stage (Stage 5): The nonlinear magnetic field is switched off when the wave packet velocity decreases to zero. A $\pi/2$ pulse is then applied to transform the state from |$\ket{0}$ to (|$(\ket{+1}+\ket{-1})/\sqrt{2}$. The linear magnetic field is then switched on. The gradient of the magnetic field is fine-tuned so that the trajectories of the two wave packets are closed after half a period.
For clarity, we use terms like ”initial separation stage” and ”Stage 1” interchangeably without ambiguity in the following sections. To simplify the discussion, only the magnetic field component in the x-direction is considered to construct the one-dimensional HP and IHP. The HP can be obtained using a linear magnetic field [2, 56]:
@@
where $B_{0}$ is a bias field along the x-direction and $\eta_{l}$ is the gradient corresponding to the linear magnetic field. Substituting Eq.(2) into the Hamiltonian Eq.(1) one have:
@@
$-\frac{\chi_{\rho}m}{2\mu_{0}}B_{0}^{2}+\hat{S}_{x}\hbar\gamma_{e}B_{0}+\hbar D\hat{S}_{z}^{2}.$ (3)
The superscript “H” denotes the HP. $\omega_{h}=(-\chi_{\rho}/\mu_{0})^{1/2}\eta_{l}$ is the frequency of the HP. The IHP can be obtained by means of a nonlinear magnetic field [59]:
@@
where $\eta_{n}$ is the gradient parameter corresponding to the nonlinear magnetic field. Note that the dimension of $\eta_{n}$ is ${\rm T/m^{2}}$. Substituting Eq.(4) into the Hamiltonian Eq.(1) one have:
@@
$-\frac{\chi_{\rho}m}{2\mu_{0}}B_{0}^{2}+\hbar D\hat{S}_{z}^{2}.$ (5)
Here $\hat{\mathbf{S}}\dotproduct\hat{\mathbf{B}}$ term has been ignored. This is because this term is zero during the enhancement and deceleration stages. It follows from Eq.(1) that the IHP is obtained when the condition $\expectationvalue{\hat{x}}\ll(2B_{0}/\eta_{n})^{1/2}$ is satisfied  [^2] When this condition is not satisfied, we get a quartic potential. This potential can also be used to create macroscopic spatial superposition [70]. The evolution of the wave packet in the quartic potential can be solved analytically and numerically [71, 72].. At this point, the Hamiltonian can be written as:
$\hat{H}_{x}^{\text{I}}=\frac{1}{2m}\hat{P}_{x}^{2}-\frac{1}{2}m\omega_{r}^{2}\hat{x}^{2}-\frac{\chi_{\rho}m}{2\mu_{0}}B_{0}^{2}+\hbar D\hat{S}_{z}^{2}.$ (6)
The superscript “I” denotes the IHP. $\omega_{r}=(-2\chi_{\rho}B_{0}\eta_{n}/\mu_{0})^{1/2}$ is the frequency of the IHP.
@@ V Wave packet evolution
If the initial state is a Gaussian shape wave packet [^6] The difference between Gaussian wave packet and Gaussian shape wave packet is that Gaussian wave packet maintains minimum uncertainty but Gaussian shape wave packets do not necessarily maintain this property [73]. (GSWP), then the evolution of the wave packet under HP and IHP can be solved analytically [74, 75, 76, 73]. In Appendix A, we give a detailed procedure for calculating the wave packet evolution using the path integral. In this section, we provide the main results. The general form of a GSWP can be written as:
@@
where $N_{0}$ is the normalization factor, $\sigma_{0}$ is the wave packet width, $x_{0}$ is the center position of the wave packet, and $a_{0}$, $b_{0}$ and $c_{0}$ are the phase related factors. The factor 1/4 in front of the parameter $a_{0}$ is set for the convenience of later calculations. A GSWP remains a GWSP after it has evolved in HP and IHP. Their general solutions can be written as:
@@
The index “I” indicates the expression of the physical quantity in the case of IHP. The expressions for the other parameters in Eq.(25) for the HP and IHP cases are given in Appendix A. The form of the solution of the wave packet in the IHP are the same as in the HP case, but with the replacement of ``$\sin"$ with ``$\sinh"$ and ``$\cos"$ with ``$\cosh"$. If the initial state is a Gaussian wave packet, the values of the parameters in Eq.(24) are:
@@
VI Fluctuation in magnetic field and wave packet contrast
Figure 5: Wave packet contrast over time when there is a small fluctuation in the linear magnetic field gradient. The red solid line denotes an initial wave packet spatial width of $\sigma_{0}=2\times 10^{-11}$ m, and the blue dashed line denotes an initial wave packet spatial width of $\sigma_{0}=3\times 10^{-11}$ m. The fluctuations in the magnetic field gradient gradually increase from left to right, for $\Delta\eta_{l}/\eta_{l}=5\times 10^{-5},1\times 10^{-4}\,\,\text{and}\,\,5\times 10^{-4}$, respectively. We set the mass $m=10^{-15}$ kg and the gradient $\eta_{l}=2507$ T/m.
In this section, we study the effect of fluctuation of the magnetic field on the contrast of the wave packets. Fluctuations in the magnetic field cause deviations from the classical position and momentum of the wave packet, preventing the two wave packets from overlapping and causing a decrease in the contrast of the wave packets. This is also known as the Humpty-Dumpty problem in the Stern-Gerlach interferometer [77].
We mark the wave packets on the two arms of the interferometer as $\psi_{\text{L}}(x,t)$ and $\psi_{\text{R}}(x,t)$. The contrast of the wave packets can be defined as:

### v2

@@ Abstract
Creating macroscopic spatial superposition states is crucial for investigating matter-wave interferometry and advancing quantum sensor technology. Currently, two potential methods exist to achieve this objective. The first involves using inverted harmonic potential (IHP) to spatially delocalize quantum states through coherent inflation [1]. The second method employs a spin-dependent force to separate two massive wave packets spatially [2]. The disadvantage of the former method is the slow initial coherent inflation, while the latter is hindered by the diamagnetism of spin-embedded nanocrystals, which suppresses spatial separation. In this study, we integrate two methods: first, we use the spin-dependent force to generate initial spatial separation, and second, we use IHP to achieve coherent inflating trajectories of the wavepackets. This approach enables the attainment of massive large spatial superposition in minimal time. For instance, a spatial superposition with a mass of $10^{-15}$ kg and a size of around 50 $\mu$m is realized in $0.1$ seconds. We also calculate the evolution of wave packets in both harmonic potential (HP) and IHP using path integral approach.
@@ I Introduction
The significant interest in massive quantum superposition states primarily arises from three aspects. The first is the exploration of the quantum-classical boundary [3]. Quantum interference, from electrons to macromolecules ($10^{-31}-10^{-23}$ kg), has been observed in contemporary experiments, demonstrating their quantum nature [4, 5, 6, 7, 8]. This leads to quests about whether quantum superposition states can also be achieved for objects of larger mass [9]. The second aspect is their utility in validating theoretical models [10]. For instance, they can be employed to test wave function collapse theories [11, 12], modified quantum mechanical frameworks [13, 14, 15], and examine the weak equivalence principle [16]. Additionally, they may reveal the quantum nature of gravity by combining two massive spatial superposition states [17, 18, 19].
The third aspect is that they can act as highly sensitive quantum sensors, detecting phenomena such as the Casimir force and dipole interactions [20, 21, 22, 23], gravitational waves [24, 25], quantum sensors for detecting accelerations, and inertial rotations [26, 27], dark matter [28], physics beyond the Standard Model [29], testing massive graviton [30], non-local gravitational interaction [31, 32], and analogue of light bending experiment in quantum gravity [33].
@@
Based on these cutting-edge techniques, numerous experimental schemes for creating macroscopic spatial superposition states have been proposed [15, 2, 48, 49, 50, 51, 1, 52, 53, 54, 55, 56, 57, 58, 23, 59]. A natural method to achieve a delocalized quantum state is to allow a quantum wave packet to evolve freely [15]. However, this delocalization process is slow. For instance, consider a silica microsphere with a mass of $10^{-15}$ kg trapped by a magnetic field at a frequency of 100 Hz [60]. Its initial wave packet spatial width is approximately $10^{-11}$ m. After 1 s of free evolution, the wave packet width becomes about $10^{-9}$ m, roughly one-thousandth of its size. To accelerate this delocalization process, an IHP can induce coherent inflation [1, 52]. This approach allows the coherent length of a $10^{-14}$ kg nanoparticle to increase to around 1 $\mu$m in 0.6 s, making the coherence length comparable to its size. Another method for creating macroscopic spatial superposition states involves utilizing spin-dependent forces, such as using diamond embedded with a NV center [2, 48, 56, 55]. Initially, a pulse is used to place the electron spin of the NV center in a superposition state, followed by applying a magnetic field to induce spatial splitting, similar to the Stern-Gerlach experiment. The advantage of this method is the ease of preparing the initial superposition state and reading out the final spin state [61]. However, the diamagnetism of the diamond suppresses the spatial separation of the wave packets when an external magnetic field is applied [54, 62, 57], limiting the direct increase of the superposition size through enhanced magnetic field gradients.
@@
The paper is organized as follows. Section II presents the specific experimental protocol and the magnetic fields required to construct the HP and IHP for the experiment. Section III provides an analytical solution for the classical trajectories of the wave packets at each experiment stage. In Section IV, we numerically calculate the classical trajectories of the wave packet without approximations for the nonlinear magnetic field and compare these results with the analytical solution. Section V discusses the evolution of the wave packet under HP and IHP using path integrals. We also examine the effect of fluctuations in the magnetic field gradient and initial position on wave packet contrast for both the HP and IHP cases in this section. Finally, we conclude our findings in Section VI.
@@ II Experimental scheme
$\hat{\mathbf{H}}=\frac{1}{2m}\hat{\mathbf{P}}^{2}-\frac{\chi_{\rho}m}{2\mu_{0}}\hat{\mathbf{B}}^{2}+\hbar\gamma_{e}\hat{\mathbf{S}}\dotproduct\hat{\mathbf{B}}+\hbar D\hat{S}_{\text{nv}}^{2},$ (1)
@@
where the first term represents the kinetic energy of the nanodiamond, with $\hat{\mathbf{P}}$ as its momentum and $m$ as its mass. The second term signifies the magnetic energy of a diamagnetic material (nanodiamond) in a magnetic field, with $\chi_{\rho}=-6.2\times 10^{-9}\,\,\text{m}^{3}/\text{kg}$ as the mass susceptibility and $\mu_{0}$ as the vacuum permeability. The third term describes the interaction between the electron spin and the magnetic field, with $\hbar$ as the reduced Planck constant, $\gamma_{e}$ as the electronic gyromagnetic ratio, $\hat{\mathbf{S}}$ as the spin operator, and $\hat{\mathbf{B}}$ as the magnetic field. The final term represents the zero-field splitting of the NV center, where $D=(2\pi)\times 2.8$ GHz and $\hat{S}_{\text{nv}}$ is the spin component operator aligned along the NV axis. Throughout this work, we assume that the nanodiamond’s rotational angular momentum is negligible. For discussions on finite angular momentum and its implications, see Ref. [62].
The experimental protocol consists of five stages:
@@
Figure 1: Schematic of the experimental scheme. The light orange quadratic curves represent harmonic or inverted harmonic potentials (HPs and IHPs). The blue and orange curves with shaded regions depict wave packets in a superposition state. The arrows inside the wave packets indicate the corresponding electron spin states (spin up or spin down). The arrows along the potential curves denote the direction of wave packet motion. The double arrows in the return stage signify the initial separation and subsequent recombination. The time axis runs from top to bottom, corresponding to Stages 1–5. The evolution times of each stage are denoted by $T_{1}$ to $T_{5}$. The gray dashed lines indicate the moments when $\pi/2$ pulses are applied.
  • Initial state:
    The system is initially prepared in the spin superposition state |$\ket{S_{x}}=(\ket{+1}+\ket{-1})/\sqrt{2}$, where |$\ket{+1}$ and |$\ket{-1}$ are eigenstates of the spin operator $\hat{S}_{x}$ with eigenvalues $+1$ and $-1$, respectively. We assume that the NV axis is aligned along the x-direction to prevent rotations due to misalignment with the magnetic field [62, 65].
  • Stage 1: Initial separation stage
    The HP is activated, causing the two wave packets to move in opposite directions, leading to a small spatial separation.
  • Stage 2: Enhancement stage
    After half a period in the HP, the wave packets reach their maximum initial separation. At this point, the HP is turned off and the IHP is activated. A $\pi/2$ pulse is then applied to transform the state |$\ket{S_{x}}=(\ket{+1}+\ket{-1})/\sqrt{2}$ into |$\ket{S_{x}}=\ket{0}$ [^1] This transformation is not strictly necessary but simplifies the equations of motion and ensures trajectory symmetry. Even without this step, the IHP can still be used to enlarge the superposition size and recombine the wave packets by fine-tuning the magnetic field. [66, 67]. The pulse duration is considered negligible. Subsequently, the wave packets rapidly separate in the IHP, achieving a large spatial superposition within a short time [^2] The evolution time depends on the desired superposition size: longer times yield greater separations..
  • Stage 3: Return stage
    The IHP is turned off, and the HP is reactivated. The wave packet velocities decrease, eventually reversing direction. The maximum spatial separation occurs at this stage.
  • Stage 4: Deceleration stage
    When the wave packet velocities reach the same magnitude (but in the opposite direction) as at the end of Stage 2333The goal of this stage is to reverse the wave packet velocity. Exact velocity matching is not strictly required but ensures trajectory symmetry. If the velocities differ, trajectory closure can still be achieved by fine-tuning the magnetic field., the HP is turned off and the IHP is turned on again. The wave packets gradually decelerate.
  • Stage 5: Recombination stage
@@
    The IHP is turned off, and a $\pi/2$ pulse is applied to restore the spin state to |$\ket{S_{x}}=(\ket{+1}+\ket{-1})/\sqrt{2}$. The HP is then activated, and the magnetic field gradient is fine-tuned to ensure that the two wave packet trajectories recombine after half a period.
This sequence is designed to achieve a large spatial superposition within a short time (approximately 0.1 s) and to ensure the recombination of the center-of-mass (CoM) trajectories, thereby restoring spin coherence [44]. Figure 1 provides an overview of these five stages. To maintain clarity, we use terms such as “Initial Separation Stage” and “Stage 1” interchangeably in subsequent discussions.
For simplicity, we consider only the x-component of the magnetic field to construct the one-dimensional HP and IHP. In realistic systems, magnetic fields must satisfy Maxwell’s equations and cannot be strictly one-dimensional. However, by applying a bias field along the x-direction and a constraining field along the y-direction, the motion of the nanodiamond can be effectively restricted to one dimension. A detailed analysis is provided in Appendix A.
The HP is generated using a linear magnetic field [2, 55]:
@@
where $B_{0}$ is the bias field along x and $\eta_{l}$ is the field gradient. Substituting Eq.(2) into the Hamiltonian Eq.(1) one have:
@@
$-\frac{\chi_{\rho}m}{2\mu_{0}}B_{0}^{2}+\hat{S}_{x}\hbar\gamma_{e}B_{0}+\hbar D\hat{S}_{\text{nv}}^{2}.$ (3)
The superscript “H” denotes the Hamiltonian associated with the HP. The corresponding frequency is given by
$\omega_{h}=\sqrt{-\frac{\chi_{\rho}}{\mu_{0}}}\eta_{l}.$ (4)
Similarly, the IHP is generated using a nonlinear magnetic field [54]:
@@
where $\eta_{n}$ is the gradient parameter characterizing the nonlinear field, with units of ${\rm T/m^{2}}$. Substituting Eq. (5) into the Hamiltonian Eq. (1), we obtain:
@@
$-\frac{\chi_{\rho}m}{2\mu_{0}}B_{0}^{2}+\hbar D\hat{S}_{\text{nv}}^{2}.$ (6)
Here, we have used the fact that $S_{x}\dotproduct B_{x}=0$ in the IHP stages, as the spin state is set to $S_{x}=\ket{0}$. From Eq. (1), the IHP is effectively realized under the condition [^4] If this condition is not met, the resulting potential takes a quartic form, which can also be employed to generate macroscopic spatial superpositions [59]. The wave packet dynamics in quartic potentials can be solved both analytically and numerically [68, 69].:
|$\absolutevalue{\expectationvalue{\hat{x}}}\ll\sqrt{\frac{2B_{0}}{\eta_{n}}}.$ (7)
Under this condition, the Hamiltonian simplifies to:
$\hat{H}_{x}^{\text{I}}=\frac{1}{2m}\hat{P}_{x}^{2}-\frac{1}{2}m\omega_{r}^{2}\hat{x}^{2}-\frac{\chi_{\rho}m}{2\mu_{0}}B_{0}^{2}+\hbar D\hat{S}_{\text{nv}}^{2}.$ (8)
The superscript “I” denotes the Hamiltonian corresponding to the IHP, with the associated frequency given by:
$\omega_{r}=\sqrt{-\frac{2\chi_{\rho}B_{0}\eta_{n}}{\mu_{0}}}.$ (9)
@@ V Wave packet evolution
If the initial state is a Gaussian shape wave packet [^8] The difference between Gaussian wave packet and Gaussian shape wave packet is that Gaussian wave packet maintains minimum uncertainty but Gaussian shape wave packets do not necessarily maintain this property [70]. (GSWP), then the evolution of the wave packet under HP and IHP can be solved analytically [71, 72, 73, 70]. In Appendix B, we give a detailed procedure for calculating the wave packet evolution using the path integral. In this section, we provide the main results. The general form of a GSWP can be written as:
@@
where $N_{0}$ is the normalization factor, $\sigma_{0}$ is the wave packet width, $x_{0}$ is the center position of the wave packet, and $a_{0}$, $b_{0}$ and $c_{0}$ are the phase related factors. The factor 1/4 in front of the parameter $a_{0}$ is set for the convenience of later calculations. A GSWP remains a GSWP after it has evolved in HP and IHP. Their general solutions can be written as:
@@
The index “I” indicates the expression of the physical quantity in the case of IHP. The expressions for the other parameters in Eq.(28) for the HP and IHP cases are given in Appendix B. The form of the solution of the wave packet in the IHP are the same as in the HP case, but with the replacement of ``$\sin"$ with ``$\sinh"$ and ``$\cos"$ with ``$\cosh"$. If the initial state is a Gaussian wave packet, the values of the parameters in Eq.(27) are:
@@
V.1 Fluctuation in magnetic field and wave packet contrast
In this section, we analyze the effect of magnetic field fluctuations on wave packet contrast. These fluctuations lead to deviations in the classical position and momentum of the wave packet, preventing the two wave packets from fully overlapping and reducing their contrast. This phenomenon is also known as the Humpty-Dumpty problem in the Stern-Gerlach interferometer [74].
We denote the wave packets in the two arms of the interferometer as $\psi_{\text{L}}(x,t)$ and $\psi_{\text{R}}(x,t)$. The wave packet contrast is defined as: